The burner on an electric stove has a diameter of .18M, Calculate the rate at which it emits energy if its temperature is 165 C. Assume it had an emissivity of 1.
Answer #1

The radiant energy Q, emitted in a time t by an object that has a Kelvin temperature T, a surface area A, and an emissivity e, is given by,

Q = e \sigma A T^4 t

where \sigma is Stefan-Boltzmann's constant with value 5.67\times 10^{-8} J/(s \cdot m^2\cdot K^4).

Then the rate of emitting energy,

\frac{dQ}{dt} = e \sigma A T^4

we have given, e = 1, radius of oven r = 0.09m, T = 165\degree C = 438 K, Then Area of the oven A = 0.0254 m^2

Then rate of energy radiated is ,

\frac{dQ}{dt} = 53.0045 J/s